# Center Biophysics and Biochemistry, (IVIC), Caracas, Venezuela

### v2

@@ Abstract
where $x_{t}$, $y_{t}$ and $z_{t}$ represent the model variables in the iteration $t$, and $\sigma$. and $\gamma$ are parameters set by the researcher before simulation (see [131] for details). In 1963 Lorenz used an analog computer where the variables of the simulated model were introduced by varying knobs that modified parameters of the simulator circuit. The analog circuit produced electrical signals that represented the results, these were printed, and before turning off the computer, Lorenz saved the values of the last set of values $x_{t+1}$. $y_{t+1}$ and $z_{t+1}$ that reintroduced into the system when it turned on the device.
As long as the system operated without interruption, the values were kept within a parallelepiped rectangle of sides $\boldsymbol{{(X_{max},Y_{max},Z_{max})}}$ (a set of its solutions is presented in Figure 1. The solution never passes through the same point twice, but evolves around that right parallelepiped forever. It seems that something attracts the system to evolve there, and even though the concept of attractor. Simply put, a attractor set A for a dynamical system is a closed subset of phase space [StateSpace2008] such that “many” (most?) of the initial conditions the system evolves towards A [150]. In the case of the Lorenz assemblage, the system solution never passes through the same place twice, a feature that coined the name strange attractor. In the Lorenz model the impossibility of reproducing or continuing from the point where it ended. This is due to the impossibility of introducing an earlier trajectory due to the imprecision of the controls of analogue systems. With digital computers, the researcher chooses an initial numerical value of his interest, this initial value is “type” always the same. At least on the same computer with the operating system and the accuracy of the processor it has, it is always handled the same and the initial (graphic) calculation is always the same. But, if the system stops and its variables are saved, follow that on Lorenz’s computer, the graph when rebooted does not follow the same trajectory: it becomes irreproducible: chaotic.
For nearly 20 years, Lorenz’s work received little attention. Only when strange attractors began to be observed in physics and then in almost every branch of science did Lorenz’s work become fundamental. Perhaps because of the shape of the Lorenz Attractor, the sensitivity to initial conditions in climate prediction was called the ”butterfly effect”. Strange attractors appeared everywhere [15]. The mathematics behind extraneous attractors can be extremely simple as follows from the set of equations (43), the implications not.
@@
As we have already said, the definition of fractal by Mandelbrot [137, 142] was almost 20 years after Lorenz’s observation [131] and introduced interest in a class of especially two-dimensional strange attractors (see Figure 12). They were not really new, the sets of Julia [111] were published 60 years earlier. Just as the development of nonparametric statistics benefited from the emergence of computers after the 1940s and their cheapening, increased accessibility in the 1950s, and cheapening with powerful and easily accessible microcomputers in the 1980s, sr advances rapidly in the processes of iterative calculation and graphing. Thus systems are developed that require iterative computation and graphing, such as systems representing chaos, iterative and fractal systems such as the Julia sets [111, 134, 124] and the beautiful and fascinating Mandelbrot set. [136, 138, 139, 143, 142, 124].
@@
Let’s look at some details about the Mandelbrot equation, it is an equation that has solutions based on a variable $Z$ and a constant $C$, both complex, as follows:
@@
where $i\!\!i=\sqrt{-1}$. The Eq. (44) is sometimes written [136] $Z_{n+1}\mapsto Z_{n}^{2}+C$. With the $Zmapping\mapsto f(Z,C)$. Where $f\in\mathbb{R}$ is a rational function $(f(Z)\in\mathbb{Q})$ of $Z$ and $C$, consider the iterated maps maps $Z_{n}=f(f(\ldots f(Z_{0})\ldots))$ of the point $Z_{0}$.
@@
$\begin{split}Z_{n}^{2}&=\left[\left(z_{re,n}^{2}-z_{im,n}^{2}\right)+2\left(z_{re,n}z_{im,n}\right)i\!\!i\right]\\ Z_{n}²&=\left[\left(z_{re,n}^{2}-z_{im,n}^{2}\right)+2\left(z_{re,n}z_{im,n}\right)i\!\!i\right]\end{split}$ (45)
@@
$\begin{split}Z_{n}^{2}&=z_{re,n}^{2}-z_{im,n}^{2}+2\left(z_{re,n}z_{im,n}\right)i\!\!i\\ Z_{n+1}^{2}&=z_{re,n}^{2}-z_{im,n}^{2}+2\left(z_{re,n}z_{im,n}\right)i\!\!i+c_{re}+c_{im}i\!\!i\\ \cdots&=\left(z_{re,n}^{2}-z_{im,n}^{2}+c_{re}\right)+\left[2\left(z_{re,n}z_{im,n}\right)+c_{im}\right]i\!\!i\\ \boldsymbol{{\text{\wasytherefore{\quad}}Z_{n+1}}}&\boldsymbol{{=\sqrt{\left\langle\left[z_{re,n}^{2}-z_{im,n}^{2}+c_{re}\right]+\left[2z_{re,n}z_{im,n}+c_{im}\right]i\!\!i\right\rangle}\quad\boldsymbol{\square}}}.\end{split}.$ (47)
@@
Mandelbrot found that the value of $\boldsymbol{{Z}}$ continued to increase or oscillate between two small “depending on the value $\boldsymbol{{C}}$)” [6]. He used a computer each value of $\boldsymbol{{C}}$ on the screen as a dot $\boldsymbol{{Z}}$ not “radial”. The result is somewhat distant a purely generic structuring process (squares, triangles, circles), each Mandelbrot graph is the solution of Eq. (44) up to $\boldsymbol{{n+1\rightarrow\aleph_{0}}}$. The resulting graph is called the Mandelbrot set. Mandelbrot continued to magnify the image, but the same structures continued to appear self-likeness). No matter when he left for $\boldsymbol{{n+1\rightarrow\aleph_{n}}}$, the same images continued to appear.
Figure 13: Mandelbrot set, Determination of $\boldsymbol{{C}}$. A window defined for complex numbers $\boldsymbol{{(x,y}}i\!\!i)$ with $\boldsymbol{{C}}$ defined in the upper right corner. Figure based on another of [6].
The procedure can be described as: create a window with complex coordinates ($x,yi\!\!i$) with one pixel per computer screen number and call it $C=x,yi\!\!i)$, the others are possible values of $Z_{n+1}\mapsto Z_{n}^{2}+C$ defined in Eq. (44). Usually, the range of the $x$ axis goes between $(\-3\leftrightarrow+3)$, and the range of $and$ values goes between $(-2\leftrightarrow+2)i\!\!i$. In this range, you can see the entire picture of the Mandelbrot set. if the values of $Z$ do not diverge is represented on site with a black dot on the screen, the result is a Mandelbrot set. Usually if $n>(2\leftrightarrow 4)$ this is greater than, it is recorded as a divergence and the recurring relationship is stopped.
@@
The introduction of the concept of chaos in climate by Lorenz [131, 130, 159], from where, after a delay, it invaded the rest of science widely, and the introduction of the concept of fractal by Mandelbrot [137, pg. 15 and Ch. 39] changed the view of the science of numerous systems, as it emerges, partially, from the previous discussion. His problem: the same as data analysis and statistics: mathematics.
@@
Using a poor method or not understanding the limitations of the method being used. In this Chapter we have presented Mandelbrot´s original definition of the concept of fractal dimension.
According to Mandelbrot [137, pg. 15 and Chap. 39]
@@
and this may not be true, co in the case of, for example, in the decimal digits of $\pi$: $N_{i}(\epsilon)\in\{0,1,\ldots,9\}$ [Sevcik2017b].In such cases, the $D_{HB}$ may be undefined, or greater, than the value of $D$ for $\left\{\log_{\epsilon}\left[N(\epsilon)\rightarrow\aleph_{0}\right]\right\}$. Under these conditionSevcikß, any empirical estimate of $D_{HB}$ will be partially uncertain. In Eq. (48, $\aleph_{0}$ is the larges value of $i$ which is $i\notin\mathbb{Z}$, the set of all integer numbers.Minimizing that uncertainty is represented by a very large set of real numbers. If $\aleph_{0}\rightarrow\infty$ the uncertainty of the decimal of $\pi$ will be represented by real numbers [Sevcik2017b]
@@
The most accurate (for even $n>2$ dimensional spaces) ways to determine $D$ were introduced by physicists [154, Theiler1990, 77, 79, 81, 154, 78]. This often involves studying the same system in Euclidean spaces of different dimensions, using Lyapunov exponents [12], and demands the use of complicated mathematics.
7.2.1 Baseless relation between fractal dimension and the Hurst´s exponent.
In the early days of the definition and “extension” of the use of the concept of fractal, Mandelbrot [137] was liberal. The association between the Hurst exponent and the fractal dimension that we discuss in this Chapter 20, we find an example.
@@
Figure 15: Properties of the ratio $d/L$ for a curve generated with simulation of Monte Carlo data [51, Y in Chapter 3 of this book.]. Panel A: of the figure shows a waveform constructed by joining with straight lines a series of uniform pseudo random numbers in the interval $(0,1)$. Panel B: is the ratio $d/L$ calculated for the corresponding point of the abscissa in the top panel. In both panels the abscissa is $N$. other details in the original publication [Sevcik1998a, Sevcik2010, Figure 2, modify] and in the text of the book.
Figure 15B shows that d/L oscillates a lot at first, but then stabilizes with $N\geqslant 50$. This confirms the above about the uselessness of Eq. (4) to determine the fractal dimension of a curve.
This is true for all waves for which ${d/L}$ is asymptotically constant after sampling ${m<\infty}$ points for ${N^{\prime}<\infty}$ the Eq. (49) also implies that the values of $D_{K}$ is (arbitrarily determined) by the choice of $N$. The boundary condition, ${N^{\prime}\rightarrow\infty}$, can be made if ${t_{max}}$ (the sampling duration) remains constant but the sampling interval ${\delta\rightarrow 0}$. I submitted Eq.(49) in my first version of my papr [Sevcik1998a, Sevcik2010] to the journal whee Katz paper appeared and was reject, more than 10 years later the published a Letter to The Editor with an “experimental proof” suggesting that there “is “something wrong” with $D_{K}$[41]. In spite of all this $D_{K}$ is still used by authors with an indecen ignorance of algebra, u ally because they know nothing else and they found it ‘pretty’ [60, 64, 61, Ripoli1999, 65, 171, 72, 71, 38, 160, Raghavendra2009, RamirezVazquez2014, 132, 86, 67].
@@
In the original version the limit is made up to $\infty$, but since $N^{\prime}\in\mathbb{C}$, it seems here more rigorous to do it up to $\aleph_{0}$. The limit (49) shows that Eq. (4) predicts that all curves are analyzed with the Eq. (4) as the discretization is improved by making $N^{\prime}$ very large or $\aleph_{0}$, we will discover that all curves are straight lines with $D_{K}=1$. It could be argued that the limit (49) has been derived assuming that $d/L$ is a constant. This is discussed in [Sevcik1998a, Sevcik2010] with the analysis of data produced with Monte Carlo simulation [51, and in Chapter 3 of this book] as shown here in Figure 15.
Also a liberal idea of Mandelbrot [137] was the idea of associating complex systems with the fractal dimension. An example of this was its association of the length of an estuary with the straight-line distance between its source and its mouth. This view inspired the fractal dimension estimated with Katz’s method [114]:
@@
This definition of Katz has a problem: only applies to short waveforms and in no case converges to fractal dimension. This is easy to see considering the limit. [Sevcik1998a, Sevcik2010].
@@
Given the inability of $D_{K}$ to predict $\Phi$, another method was developed to calculate the fractal dimension [Sevcik1998a, Sevcik2010], this method describes what is now called Sevcik’s fractal dimension in GogScholar [5] refers 198 () citations of the use of Sevcik´s fractal dimension [Sevcik1998a, Sevcik2010], using this algorithm.The [Sevcik1998a, Sevcik2010] algorithm is a double linear transformation [Eqs. (8) – (15)].
As said $D_{S}$ is calculated in a linear transformation of the space where the waveform exist, thus both spaces have the same metric properties, whih includes $D_{HB}$, $\Phi$ and $D_{S}$ [20, Sevcik1998a, Sevcik2010] if $\boldsymbol{{N\longrightarrow\infty}}$. Since the topology of a metric space does not change under a linear transformation, it is convenient to transform a wave into another normalized space, where all axes are equal. I proposed the use of two linear transformations that map the original wave function into another embedded in an equivalent metric space [please see [20] for a very readable discussion of metric spaces]. In recent years, this way of calculating the fractal dimension of a wave has been called the Sevcik fractal dimension [RodriguezHernandez2022], which here we call $D_{S}$. This determines that the original function and its transform without plane curves into a space with coordinates with units.
Thus it is not surprising that, as shown in Subs-subsection 3.2.2, $\boldsymbol{{D_{S}}}$ converges to $\boldsymbol{{\Phi}}$ as $\boldsymbol{{N\rightarrow\infty}}$.
@@
Fortunately $\approx 10^{1}4$ decimals of $\pi$ are current
Part I title
ly known [3], and $10^{9}$ are easy to generate on a personal computer [23, Sevcik2017b]. Thus many sequences of series of random uniformly distributed each with a probability $\frac{1}{10}$ decimal digits $N$ and compared with the same numbers of $\pi$ decimals [Sevcik2017b]. The same procedure may be used to compare an observed series and compare ot with a random sequence of known properties.
@@
Although it is possible to prove for fractal functions such as Koch snowflake [Koch1904, Sevcik1998a, Sevcik2010] that $\underset{N\rightarrow\aleph_{0}}{\lim}D_{S}=D_{HB}$, for more complex functions this can be difficult [Sevcik1998a, Sevcik2010, Sevcik2017b, RodriguezHernandez2022] ka demonstration may not be easy and require Monte Carlo simulations [Shi2012, 151, Sevcik1998a, Sevcik2010, Sevcik2017b, RodriguezHernandez2022].
Intuitive estimations of $\Phi$ are dangerous. Even a solid mathematician as Mandelbrot devised more or less intuitive estimates for rivers using the length of the river and the straight line distance between the river fountain to the river end [137], the Katz $D_{K}$ fractal dimension is a generalization Mandelbrot§river $Psi$. Even if the fractal dimension defined by Mandelbrot is the ratio of two “natural” river, perhaps better for long rives, but $D_{K}$ gets worse as the number of discretized waveform points grows [114, Sevcik1998a, Sevcik2010].
@@
The term fractal is sometimes used as as manner to provide a ‘deep scientific’ meaning, to propositions lacking any scientific sport. This seems the case this seems the case of a theory unifying gravity and quantum physics [146, 147, among several YouTube videos and a LinkedIn comunication] or with fancy sounding theories to synthesize drugs [14].
@@
Authors make mistakes and reviewers not always understand the concept of fractal dimension. Between the non physicists, fractal analysis became fashionable in this century. Unfortunately, too many authors (and their reviewers in journal) do not pay attention to the mathematics of the fractal analysis methods, thus the Katz and Higuchi methods have gained undue popularity, specially medical, psychological, psychiatric an low level technique papers, there are too many examples to cithem precisely here but many examples can be found with GoogleScholar (https://bitly.ws/347Cn) or with sites such as Academia.edu (https://bitly.ws/347EA) (among other scientific paper browsers) intermixed with valuable papers.
Here we are going to consider a problem (excuse the neologism) that we will call: the “curve nicety”: a principle that demands that, above any other principle, the curves and data that we enter in a work must look ‘pretty’. Of course, if they are also true better than better. Some cases are in our Chapter 10. While Eq. (49) [Sevcik1998a, 41, Sevcik2010] demonstrates beyond doubt that the call “fractal dimension of Katz” [114] is different from $1$ only if the starting points of any waveform are considered and is equal to 1 for any waveform other than a straight line when $N\rightarrow\aleph_{0}$ or $N\rightarrow\infty$ and $\underset{N\rightarrow\infty}{\lim}D_{HB}\neq 1.0$. The problem is that Eq. (4) does not have an “optimal” to estimate $D_{K}$ Eq. (49), always underestimates it even if you underestimate it more with large $N$. However a number of authors base their choice of a method of assigning fractal dimension to their curves choose the one that makes their data look “prettier” and thus preserve meaningless methods [60, 64, 61, Ripoli1999, 65, 171, 72, 71, 38, 160, Raghavendra2009, RamirezVazquez2014, 132, 86, 67, 2, 55, Klonowski2005, Spacic2011, Spasic2011a, 90, 1, 49, 125, 126, Shamsi2021, 9].

### v3

@@ Abstract
where $x_{t}$ , $y_{t}$ and $z_{t}$ represent the model variables in the iteration $t$, and $\sigma$. and $\gamma$ are parameters set by the researcher before simulation (see [133] for details). In 1963 Lorenz used an analog computer where the variables of the simulated model were introduced by varying knobs that modified parameters of the simulator circuit. The analog circuit produced electrical signals that represented the results, these were printed, and before turning off the computer, Lorenz saved the values of the last set $\{x_{t+1},y_{t+1},z_{t+1}\}$ that he reintroduced into the system when he turned on the device the next time.
As long as the system operated without interruptions, the values stayed within a parallelepiped rectangle of sides $\boldsymbol{{\{X_{max},Y_{max},Z_{max}\}}}$ (a set of its solutions is presented in Figure LABEL:F:Lorenz). The solution never passes through the same point twice, but evolves around that right parallelepiped forever. It seems as if something attracts the system to evolve there, thus the concept of attractor was coined to describe situations like this. Simply put, a attractor set A for a dynamical system is a closed subset of phase space [StateSpace2008] such that “many” (most?) of the initial conditions the system evolves towards A [152]. In the case of the Lorenz ensemble, the system solution never passes through the same place twice, a feature that coined the name strange attractor. In the Lorenz model the impossibility of reproducing or continuing from the point where it ended. This is due to the impossibility of introducing an earlier trajectory due to the imprecision of the controls of analogue systems. With digital computers, the researcher chooses an initial numerical value of his interest, this initial value is “type” always the same. At least on the same computer with the operating system and the accuracy of the processor it has, it is always handled the same and the initial (graphic) calculation is always the same. But, if the system stops and its variables are saved, behave the same as the analog Lorenz’s computer, the graph when rebooted does not follow the same trajectory: it becomes irreproducible: chaotic.
For nearly 20 years, Lorenz’s work received little attention. Only when strange attractors began to be observed in physics and then in almost every branch of science did Lorenz’s work became fundamental. Perhaps because of the shape of the Lorenz Attractor, the sensitivity to initial conditions in climate prediction was called the “butterfly effect”. Strange attractors appeared everywhere [15]. The mathematics behind strange attractors can be extremely simple as follows from the set of equations (46), but the implications deep.
@@
As we have already said, the definition of fractal by Mandelbrot [139, 144] occurred almost 20 years after Lorenz’s work [133] and introduced interest in a class of, especially, two-dimensional strange attractors (see Figure 12). They were not really new, the sets of Julia [113] were published 60 years earlier. Just as the development of nonparametric statistics benefited from the emergence of computers after the 1940s, and their cheapening increased accessibility in the 1950s and cheapening with powerful and easily accessible microcomputers in the 1980s. As consequence, rapid advances occurred in processioning of iterative calculation and graphing. Thus systems that require iterative computation and graphing were developed. Such as systems, representing chaos, iterative and fractal systems such as the Julia sets [113, 136, 126] and the beautiful and fascinating Mandelbrot set. [138, 140, 141, 145, 144, 126].
@@
Let’s look at some details about the Mandelbrot equation, it is an equation that has solutions based on a variable $Z\in\mathbb{C}$ and a constant $C\in\mathbb{C}$, as follows:
@@
where $i\!\!i=\sqrt{-1}$. Eq. (47) is sometimes written [138] $Z_{n+1}\mapsto Z_{n}^{2}+C$. With the $Zmapping\mapsto f(Z,C)$. Where $f\in\mathbb{R}$ is an irrational function ($\mathbb{Q}$) so that $(f(Z)\in\mathbb{Q})$ of $Z$ and $C$, consider the iterated maps $Z_{n}=f(f(\ldots f(Z_{0})\ldots))$ of the point $Z_{0}$.
@@
$\begin{split}Z_{n}²&=\left[\left(z_{re,n}^{2}-z_{im,n}^{2}\right)+2\left(z_{re,n}z_{im,n}\right)i\!\!i\right]\end{split}$ (48)
@@
$\begin{split}Z_{n}^{2}&=z_{re,n}^{2}-z_{im,n}^{2}+2\left(z_{re,n}z_{im,n}\right)i\!\!i\\ Z_{n+1}^{2}&=z_{re,n}^{2}-z_{im,n}^{2}+2\left(z_{re,n}z_{im,n}\right)i\!\!i+c_{re}+c_{im}i\!\!i\\ \cdots&=\left[z_{re,n}^{2}-z_{im,n}^{2}+c_{re}\right]+\left[2\left(z_{re,n}z_{im,n}\right)+c_{im}\right]i\!\!i\\ \boldsymbol{{\text{\wasytherefore{\quad}}Z_{n+1}}}&\boldsymbol{{=\sqrt{\left[z_{re,n}^{2}-z_{im,n}^{2}+c_{re}\right]+\left[2z_{re,n}z_{im,n}+c_{im}\right]i\!\!i}\quad\boldsymbol{\square}}}.\end{split}.$ (50)
@@
Mandelbrot found that the value of $\boldsymbol{{Z}}$ continued to increase or oscillate between two smallvalues “depending on $\boldsymbol{{C}}$)” [5]. He used a computer to calculate each value of $\boldsymbol{{C}}$ and plotted it on the screen as a not “radial” dot $\boldsymbol{{Z}}$. The result is somewhat difficult to describe, a purely generic structuring process (squares, triangles, circles), each Mandelbrot graph is the solution of Eq. (47) up to $\boldsymbol{{n+1\rightarrow\aleph_{0}}}$. The resulting graph is called the Mandelbrot set. Mandelbrot continued to magnify the image, but the same structures continued to appear (self-likeness). No matter when he left, for $\boldsymbol{{n+1\rightarrow\aleph_{n}}}$, the same images continued to appear.
Figure 13: Mandelbrot set. Determination of $\boldsymbol{{C}}$. A window defined for complex numbers $\boldsymbol{{(x,y}}i\!\!i)$ with $\boldsymbol{{C}}$ defined in the upper right corner. Figure based on another of [5].
The procedure can be described as: create a window with complex coordinates ($x,yi\!\!i$) with one pixel per computer screen number and call it $C=x,yi\!\!i)$, the others are possible values of $Z_{n+1}\mapsto Z_{n}^{2}+C$ defined in Eq. (47). Usually, the range of the $x$ axis goes between $(-3\leftrightarrow+3)$, and the range of $y$ values goes between $(-2\leftrightarrow+2)i\!\!i$. In this range, you can see the entire picture of the Mandelbrot set. if the values of $Z$ do not diverge is represented on site with a black dot on the screen, the result is a Mandelbrot set. Usually if $n>(2\leftrightarrow 4)$ this is greater than, it is recorded as a divergence and the recurring relationship is stopped.
@@
The introduction of the concept of chaos in climate by Lorenz [133, 132, 161], from where, after a delay, it invaded the rest of science widely, and the introduction of the concept of fractal by Mandelbrot [139, pg. 15 and Ch. 39] changed the view of the science of numerous systems, as it emerges, partially, from the previous discussion. His problem: the same as data analysis and statistics: mathematics.
@@
A fundamental problem in fractal analysis is using a poor method or not understanding the limitations of the method being used. In this paper we have presented Mandelbrot´s original definition of the concept of fractal dimension.
According to Mandelbrot [139, pg. 15 and Chap. 39]
@@
but this may not be true, in the case of, for example, the decimal digits of $\pi$: $N_{i}(\epsilon)\in\{0,1,\ldots,9\}$ [Sevcik2017b]. In such cases, $D_{HB}$ may be undefined, or greater, than the value of $D$ for $\left\{\log_{\epsilon}\left[N(\epsilon)\rightarrow\aleph_{0}\right]\right\}$. Under these condition Sevcik´s $D_{S}$, or any empirical estimate of $D_{HB}$ will be partially uncertain. In Eq. (51), $\aleph_{0}$ is the larges value of $i$ which is $(i\in\mathbb{Z})$, the set of all integer numbers. Minimizing that uncertainty is represented by a very large set of numbers. If $\aleph_{0}\rightarrow\infty$ the uncertainty of the decimal of $\pi$ will be represented by real numbers [Sevcik2017b]
@@
The most accurate (for even $n>2$ dimensional spaces) ways to determine $D$ were introduced by physicists [156, Theiler1990, 79, 81, 83, 156, 80]. This often involves studying the same system in Euclidean spaces of different dimensions, using Lyapunov exponents [10], and demands the use of complicated mathematics.
7.2.1 Baseless relation between fractal dimension and the Hurst´s coefficient ($H$).
In the early days of the definition and “extension” of the concept of fractal, Mandelbrot [139] was liberal. The association between the Hurst exponent and the fractal dimension that we discuss in this here, we find an example.
@@
Figure 15: Properties of the ratio $\boldsymbol{{d/L}}$ for a curve generated with Monte Carlo simulation [52]. Panel A: Shows a waveform constructed by joining with straight lines a series of uniform pseudo random numbers in the interval $(0,1)$. Panel B: Ratio $d/L$ calculated for the corresponding point of the abscissa in the top panel. In both panels the abscissa is $N$. other details in the original publication [Sevcik1998a, Sevcik2010, Figure 2] and in the text of the this paper.
Figure 15B shows that $d/L$ oscillates a lot at first, but then stabilizes with $N\geqslant 50$. This confirms the uselessness of [Eq. (4)] to determine the fractal dimension of a curve. This is true for all waves for which ${d/L}$ is asymptotically constant after sampling ${m<\infty}$ points for ${N^{\prime}<\infty}$ Eq. (52) also implies that the values of $D_{K}$ is arbitrarily determined by the choice of $N$. The boundary condition, ${N^{\prime}\rightarrow\infty}$, can be made if ${t_{max}}$ (the sampling duration) remains constant but the sampling interval ${\delta\rightarrow 0}$. I submitted Eq.(52) in the first version of my papr [Sevcik1998a, Sevcik2010] to the journal whee Katz paper appeared and was rejected, more than 10 years later the published a Letter to The Editor with an “experimental proof” suggesting that there is “something wrong” with $D_{K}$ [42]. In spite of all this $D_{K}$ is still used by authors with ignorance of algebra, usally because they know nothing else and they found it ‘pretty’ [61, 66, 63, Ripoli1999, 67, 173, 74, 73, 39, 162, Raghavendra2009, RamirezVazquez2014, 134, 88, 69].
@@
In the original version [Sevcik1998a, Sevcik2010] the limit is made up to $\infty$, but since $N^{\prime}\in\mathbb{Z}$, it seems here more rigorous to do it up to $\aleph_{0}$. The limit in Eq. (52) shows that Eq. (4) predicts that all curves are analyzed with the Eq. (4) as the discretization is improved by making $N^{\prime}$ very large or $\aleph_{0}$, we will discover that all curves are straight lines with $D_{K}=1$. It could be argued that the limit in Eq. (52) has been derived assuming that $d/L$ is a constant. This is discussed in [Sevcik1998a, Sevcik2010] with the analysis of data produced with Monte Carlo simulation [52] as shown in Figure 15.
Also a ‘liberal’ idea of Mandelbrot [139] was associating complex systems with the fractal dimension. An example of this was its association of the length of a river with a straight line distance between its source and its mouth. This view inspired the fractal dimension estimated with Katz’s method [116]:
@@
This definition of Katz has a problem: only applies to short waveforms and in no case converges to $\boldsymbol{{\Phi}}$. This is easy to see considering the limit in Eq.(52) [Sevcik1998a, Sevcik2010].
@@
[…]
As said $D_{S}$ is calculated in a linear transformation of the space where the waveform exist, thus both spaces have the same metric properties, whih includes $D_{HB}$, $\Phi$ and $D_{S}$ [20, Sevcik1998a, Sevcik2010] if $\boldsymbol{{N\longrightarrow\infty}}$. Since the topology of a metric space does not change under a linear transformation, it is convenient to transform a wave into another normalized space, where all axes are equal. I proposed the use of two linear transformations that map the original wave function into another embedded in an equivalent metric space [please see [20] for a very readable discussion of metric spaces]. In recent years, this way of calculating the fractal dimension of a wave has been called the Sevcik fractal dimension [RodriguezHernandez2022], which here we call $D_{S}$. This determines that the original function and is transformed into anther into a space with equal coordinates and with units. Thus it is not surprising that, as shown in Subs-subsection 3.2.2, $\boldsymbol{{D_{S}}}$ converges to $\boldsymbol{{\Phi}}$ as $\boldsymbol{{N\rightarrow\infty}}$.
@@
As of July 2024, $Pi$ has been calculated to $2.02\cdot 10^{14}$ digits [13], and $10^{9}$ are easy to generate on a personal computer [23, Sevcik2017b]. Thus many sequences of series of random uniformly distributed each with a probability $\frac{1}{10}$ decimal digits $N$ and compared with the same numbers of $\pi$ of decimals [Sevcik2017b]. The same procedure may be used to compare an observed series and compare ot with a random sequence of known properties.
@@
Although it is possible to prove for fractal functions such as Koch snowflake [Koch1904, Sevcik1998a, Sevcik2010] that $\underset{N\rightarrow\aleph_{0}}{\lim}D_{S}=D_{HB}$, for more complex functions this can be difficult [Sevcik1998a, Sevcik2010, Sevcik2017b, RodriguezHernandez2022] ka demonstration may not be easy and require relianle Monte Carlo simulations [Shi2012, 153, Sevcik1998a, Sevcik2010, Sevcik2017b, RodriguezHernandez2022].
Intuitive estimations of $\Phi$ are dangerous. Even a solid mathematician as Mandelbrot devised more or less intuitive estimates for rivers using the length of the river and the straight line distance between the river fountain to the river end [139], the Katz $D_{K}$ fractal dimension is a generalization Mandelbrot´s river $\Psi$. Even if the fractal dimension defined by Mandelbrot is the ratio of two “natural” river, perhaps better for long rives, but $D_{K}$ gets worse as the number of discretized waveform points grows [116, Sevcik1998a, Sevcik2010].
@@
The term fractal is sometimes used as as manner to provide a ‘deep scientific’ meaning, to propositions lacking any scientific sport. This seems the case this seems the case of a theory unifying gravity and quantum physics [148, 149, among several YouTube videos and a LinkedIn comunication] or with fancy sounding theories to synthesize drugs [14].
@@
Authors make mistakes and reviewers not always understand the concept of fractal dimension. Between the non physicists, fractal analysis became fashionable in this century. Unfortunately, too many authors (and their reviewers in journals) do not pay attention to the mathematics of the fractal analysis methods, thus the Katz and Higuchi methods have gained undue popularity, specially in medical, psychological, psychiatric an low level technique papers, there are too many examples to cithem precisely here but many examples can be found with GoogleScholar (https://scholar.google.com) or with sites such as Academia.edu (https://www.academia.edu) (among other scientific paper browsers) intermixed with valuable papers.
Another problem is (excuse the neologism) that we will call: the “curve nicety”: a principle that demands that, above any other principle, the curves and data that we enter in a work must look ‘pretty’. Of course, if they are also true better than better. While Eq. (52) [Sevcik1998a, 42, Sevcik2010] demonstrates beyond doubt that the called “fractal dimension of Katz” [116] is different from $1$ only if the starting points of any waveform are considered and is equal to 1 for any waveform other than a straight line when $N\rightarrow\aleph_{0}$ or $N\rightarrow\infty$ and $\underset{N\rightarrow\infty}{\lim}D_{HB}\neq 1.0$. The problem is that Eq. (4) does not have an “optimal” to estimate $D_{K}$ Eq. (52), always underestimates $\Phi$ even if you underestimate it more with large $N$. However a number of authors base their choice of a method of assigning fractal dimension to their curves choose the one that makes their data look “prettier” and thus preserve meaningless methods [61, 66, 63, Ripoli1999, 67, 173, 74, 73, 39, 162, Raghavendra2009, RamirezVazquez2014, 134, 88, 69, 2, 56, Klonowski2005, Spacic2011, Spasic2011a, 92, 1, 50, 127, 128, Shamsi2021, 8].